OpenStudy (anonymous):

What is the solution to the following system of equations? y^2+x^2=53 y-x=5 a)(-10,-5) b)(-7,-2) c)(7,2) d)(5,10)

OpenStudy (gudden):

y - x = 5 => y = 5 + x plug y == 5+x, in y^2+x^2=53

OpenStudy (gudden):

@15Twilight ???

OpenStudy (anonymous):

ok but then I get (5+x)^2+x^2=53 what do I do next

OpenStudy (gudden):

open the (5 +x)^2, using two degree binomial expansion ( in simple words (a+b)^2 = a^2 + b^2 +2ab )

OpenStudy (anonymous):

thank you @gudden